### v2

# Aligning Language Models with Real-time Knowledge Editing

Abstract
Knowledge editing aims to modify outdated knowledge in large language models (LLMs) efficiently while retaining their original capabilities. Mainstream benchmarks for knowledge editing are predominantly static and fail to keep in pace with the evolving real-world knowledge. In this work, we introduce CRAFT, an ever-evolving real-world benchmark for knowledge editing. It features well-designed paired edits for composite reasoning, and evaluates models on alias portability as well as temporal and common-sense locality, making it a challenging knowledge editing benchmark on which previous knowledge editing methods hardly achieve balanced performance. Towards flexible real-time editing, we propose KEDAS, a novel paradigm of knowledge editing alignment featuring diverse edit augmentation and self-adaptive post-alignment inference, which exhibits significant performance gain on CRAFT compared to previous methods. All of our code and data are available at https://anonymous.4open.science/r/CRAFT-KEDAS.
Aligning Language Models with Real-time Knowledge Editing
Chenming Tang   Yutong Yang   Kexue Wang   Yunfang Wuthanks: Corresponding author. National Key Laboratory for Multimedia Information Processing, Peking University MOE Key Laboratory of Computational Linguistics, Peking University School of Computer Science, Peking University {tangchenming, 2300013219}@stu.pku.edu.cn {yytpku, wuyf}@pku.edu.cn
$\dagger$$\dagger$}\$footnotetext: These authors contributed equally.
1 Introduction
Large language models (LLMs) Llama Team (2024); Gemini Team (2025) have been the core of modern natural language processing (NLP). Once pre-trained, knowledge is injected into LLMs and becomes their static internal capability Petroni et al. (2019). As time goes by, some knowledge in LLMs inevitably becomes incorrect or out of date, and it is vital to update the outdated knowledge. However, retraining a model from scratch is highly costly, especially for downstream users. To this end, the techniques of knowledge editing have been developed Zhang et al. (2024), aimed at editing specific knowledge in LLMs efficiently without heavy re-training.
There have been a large number of benchmarks to evaluate the effectiveness of knowledge editing methods. Nevertheless, most traditional knowledge editing benchmarks like ZsRE Levy et al. (2017), MQuAKE Zhong et al. (2023) and RippleEdits Cohen et al. (2024a) are static and not real-time. Once the datasets are constructed and released, they remain fixed and can not be updated anymore, which is far from real-world knowledge editing applications. Recently, some work is devoted to building real-time datasets. For example, WikiBigEdit Thede et al. (2025) collects real-world changes in Wikipedia, but it requires processing massive Wiki data (usually hundreds of gigabytes) from the Internet, which is costly and inconvenient in practice. Meanwhile, the automatically collected dataset lacks necessary filtering and thus suffers from severe sparsity (for example, only 3.6% of the collected data have an entry of portability evaluation). EvoWiki Tang et al. (2025) is another real-time dataset also based on Wikipedia, but it focuses on retrieval-augmented generation (RAG) and continual learning. Overall, there still lacks a high-quality, easily-updated benchmark for real-world and real-time knowledge editing.
Benchmark Real- time Real- world Reasoning Portability Composite Reasoning Portability
ZsRE
MQuAKE-T
MQuAKE-CF
RippleEdits
WikiBigEdit
EvoWiki $\triangle$
CRAFT (Ours)
Table 1: Comparison of knowledge editing benchmarks. Composite reasoning portability evaluates a model’s ability to integrate multiple distinct edits into a single reasoning query, whereas reasoning portability tests reasoning chains derived from a single edit.
In this work, we introduce CRAFT (Chinese Real-time statistics And Finance knowledge ediTing benchmark), an ever-evolving real-world benchmark for knowledge editing in Chinese. It leverages publicly available official data that are continuously updated, ensuring both temporal freshness and real-world applications. Furthermore, it is organized in well-designed paired edits, each serving as a composite reasoning test to evaluate models’ ability to integrate multiple related factual updates. Moreover, CRAFT supports evaluations on alias portability and temporal and common-sense locality, providing a comprehensive assessment of model adaptability and robustness under dynamic knowledge changes. The differences between CRAFT and previous datasets are summarized in Table 1.
We assess the exposure rate of CRAFT and existing benchmarks with five different LLMs, either open-source or closed-source. The results demonstrate that a large portion of knowledge in existing benchmarks has been known to LLMs while the exposure rate of CRAFT is nearly zero on all the five models. This reveals the disadvantages of static datasets and validates the real-time property of CRAFT, which makes it challenging to LLMs.
We evaluate a suite of representative knowledge editing methods on CRAFT and demonstrate that current approaches exhibit inherent limitations. For example, parameter-based approaches like ROME Meng et al. (2022) and WISE Wang et al. (2024a) suffer from gradual model degradation and struggle to achieve successful editing in the setting of sequential editing. Simple retrieval-based approaches like IKE Zheng et al. (2023) and EREN Chen et al. (2024b) fail to achieve consistent balanced performance due to lack of alignment. LTE Jiang et al. (2024), which aligns LLMs with knowledge editing via post-training, shows weak locality due to overfitting on irrelevant queries.
[…]
Our contributions are as follows:
  • We introduce CRAFT, a high-quality, easily-updated and fully open-source benchmark for real-time knowledge editing.
  • We evaluate existing knowledge editing methods on CRAFT and reveal the limitations of these methods in real-time knowledge editing.
  • We propose KEDAS, a novel paradigm of aligning LLMs with real-time knowledge editing featuring diverse edit augmentation and self-adaptive post-alignment inference, enabling dynamic routing between pre- and post-alignment LLMs, which significantly outperforms existing methods on CRAFT.
2 Preliminary
2.1 Task Formulation
An LLM system $f$ can be regarded as a function $f:\mathcal{Q}\mapsto\mathcal{A}$ mapping an query $q\in\mathcal{Q}$ to its output answer $a=f(q)\in\mathcal{A}$.
Knowledge editing aims to change the behavior of the LLM after modifying some knowledge. Edits of knowledge are usually in the form of query-answer (QA) pairs:
$\mathcal{E}=\{e^{t}\}^{N}_{t=1}=\{(q^{t}_{e},a^{t}_{e})\}^{N}_{t=1},$ (1)
where $q^{t}_{e}$ is an input query triggering knowledge (e.g., The current US president is), $a^{t}_{e}$ is the corresponding target of editing (e.g., Donald Trump), $t$ is the index of each edit, and $N$ denotes the total number of edits.
To assess the efficacy of editing, the post-edit LLM $f^{*}$ is evaluated via the following criteria Zhang et al. (2024):
Edit success
measures the accuracy of editing, requiring $f^{*}$ to correctly recall the edits:
$\mathbb{E}_{(q_{e},a_{e})\in\mathcal{E}}\ \mathds{1}\{f^{*}(q_{e})=a_{e}\}.$ (2)
Locality
measures the precision of editing, requiring $f^{*}$ not to change its behavior out of the scope of edits:
$\mathbb{E}_{(q_{l},a_{l})\in\mathcal{E}_{l}}\ \mathds{1}\{f^{*}(q_{l})=f(q_{l})\},$ (3)
where $\mathcal{E}_{l}$ are QA pairs of unrelated queries. Note that the target of locality is unchanged answer rather than exactly the ground truth answer.
Portability
measures how well edits transfer to related queries, requiring $f^{*}$ to correctly answer such queries:
$\mathbb{E}_{(q_{p},a_{p})\in\mathcal{E}_{p}}\ \mathds{1}\{f^{*}(q_{p})=a_{p}\},$ (4)
where $\mathcal{E}_{p}$ are QA pairs related to existing edits.
2.2 Settings of Knowledge Editing
Single Editing
In single editing, each edit is evaluated directly after it is applied to the original model $f$:
$f^{t}=\mathbf{Edit}(f,q^{t}_{e},a^{t}_{e}),\quad 1\leq t\leq N.$ (5)
This traditional setting is widely employed by earlier work. It is far from real-world applications, since only one edit can be applied each time.
Sequential Editing
In sequential editing, edits are applied step by step (let $f^{0}$ denote $f$ for consistency):
$f^{t}=\mathbf{Edit}(f^{t-1},q^{t}_{e},a^{t}_{e}),\quad 1\leq t\leq N.$ (6)
The final model $f^{N}$ after applying all $N$ edits is evaluated. This setting is frequently used to evaluate the lifelong performance and scalability of knowledge editing. We focus on this setting in this work because it is closer to practical applications, especially the real-time knowledge editing setting.
Knowledge Editing Alignment
LTE Jiang et al. (2024) relates knowledge editing to LLM alignment, aligning LLMs with ever-changing real-time knowledge edits. During alignment, LLMs’ capabilities of knowledge updating are enlightened by an editing prompt:
    [Updated Information] {edit}
    [Query] {query}
The training data consist of in-scope and out-of-scope queries with or without the specific edit to improve edit efficacy while keeping locality.
In the inference phase, LLMs are required to conduct on-the-fly and streaming knowledge editing by retrieving relevant edits to the query from the stored memory.
Dataset GPT-3.5 Mistral-7B Llama-3.1-8B DeepSeek-chat Doubao-1.5-pro-32k Size
OpenAI (2023) Jiang et al. (2023) Llama Team (2024) DeepSeek-AI (2025) ByteDance (2025)
ZsRE Levy et al. (2017) 43.98% 23.48% 22.45% 52.47% 26.77% 6,000
MQuAKE Zhong et al. (2023) 35.46% 76.21% 46.11% 42.58% 76.09% 5,604
EvoWiki_evolved Tang et al. (2025) 8.07% 16.22% 6.43% 10.24% 5.08% 1,338
EvoWiki_stable Tang et al. (2025) 28.45% 42.37% 25.90% 39.76% 12.32% 1,494
EvoWiki_uncharted Tang et al. (2025) 8.19% 22.62% 7.39% 16.02% 3.43% 1,136
CRAFT-Statistics (Ours) 0.01% 0.00% 0.30% 1.80% 0.99% 4,468
CRAFT-Finance (Ours) 0.00% 0.00% 0.14% 0.42% 0.00% 7,800
Table 2: Exposure rate (%) of different LLMs to benchmark datasets. Model release years are indicated in citations. CRAFT exhibits minimal exposure due to real-time data collection and domain freshness.
3 The CRAFT Dataset
We introduce CRAFT, a real-time, real-world benchmark designed to evaluate knowledge editing under dynamically evolving factual contexts. Unlike static benchmarks that rely on outdated or widely exposed Wikipedia facts, CRAFT continuously updates with publicly available official data sources, ensuring fairness and real-world relevance.
3.1 Dataset Overview
CRAFT consists of two complementary subsets: CRAFT-Statistics and CRAFT-Finance, representing stable and highly correlated factual domains that evolve periodically. Example instances from the two subsets are shown in Appendix E and F respectively.
CRAFT-Statistics
We collect monthly statistical reports from the official National Bureau of Statistics of China 222 https://data.stats.gov.cn/, covering July 2024 to June 2025 in the current version used in this work. The dataset spans four major domains—finance, fiscal affairs, telecommunications, and transportation—containing 221 indicators and 4,468 data points. Our data collection pipeline, leveraging the cn-stats API 333 https://github.com/songjian/cnstats, supports automatic temporal updates, enabling effortless retrieval of the most recent monthly data.
CRAFT-Finance
CRAFT-Finance includes annual financial statements of 390 publicly listed Chinese companies (2023–2024), sourced from Eastmoney Financial Database 444 https://data.eastmoney.com/ via the AKShare API 555 https://github.com/akfamily/akshare. It features financial indicators such as debt ratio, shareholder equity, and operating cash flow, supporting yearly updates to reflect the latest financial disclosures.
3.2 Evaluation Metrics
To comprehensively assess model adaptability, besides the conventional edit success (ES) as described in Section 2.1, CRAFT provides four evaluation dimensions:
[…]
We propose composite reasoning portability, a new evaluation designed specifically for CRAFT. Unlike traditional multi-hop reasoning portability, which constructs reasoning chains from a single edit, composite portability evaluates a model’s ability to integrate multiple distinct edits into a single reasoning query.
In CRAFT, all data are organized in paired format: each test instance consists of two edits and one corresponding composite reasoning query. Formally, given a pair of edits $\{e_{1},e_{2}\}$, the composite portability query $q_{p}$ requires reasoning over both edits simultaneously. For example, consider a CRAFT-Finance instance with a pair of edits:
$e_{1}$ $:\text{2024 Company A's total assets = 2M}$
$e_{2}$ $:\text{2024 Company A's total liabilities = 1M}$
The corresponding composite reasoning query and answer are:
$q_{p}$ $=\text{2024 Company A's debt ratio}$
$a_{p}$ $=0.5$
CRAFT-Finance includes ten types of such combinations, demonstrating the generality and practical relevance of composite reasoning portability. For CRAFT-Statistics, composite portability is constructed by pairing two edits of the same indicator from adjacent years.
[…]
We propose alias portability to evaluate a model’s robustness to synonymous indicator names by replacing key terms with domain-specific aliases constructed from a manually curated alias list.
[…]
We propose temporal locality to measure a model’s sensitivity to temporal shifts by querying facts from neighboring time periods.
[…]
3.3 Dataset Statistics
Table 2 compares the exposure ratio of CRAFT and existing benchmarks across five representative LLMs. Due to its temporal freshness and domain specificity, CRAFT exhibits extremely low exposure rates across all the five LLMs, mitigating data contamination and ensuring fair evaluation. Meanwhile, for other datasets, a great portion of the knowledge has been known to LLMs, making them less challenging and less fair in practice.
Dataset Edit $\mathbf{P}_{\text{reasoning}}$ $\mathbf{P}_{\text{alias}}$ $\mathbf{L}_{\text{temporal}}$ $\mathbf{L}_{\text{common}}$
CRAFT- Statistics 4,468 2,234 4,468 4,468 4,468
CRAFT- Finance 7,800 3,900 7,800 7,800 7,800
Table 3: Number of instances per evaluation type in CRAFT. Composite reasoning uses edit pairs; other metrics are per edit.
Table 3 shows instance counts in CRAFT sub-datasets. Composite reasoning portability is evaluated per edit pair, and other metrics per edit. For experiments, we split CRAFT into training and test sets. The test set contains 500 instances from the two subsets each (1,000 instances in total) and the training set contains the remaining instances.
3.4 Data Summary
CRAFT offers a scalable and continuously evolving benchmark reflecting real-world knowledge updates. Its dynamic nature, factual authenticity, and comprehensive reasoning evaluation make it a robust foundation for future research on reliable and interpretable knowledge editing. Anyone who wants to construct CRAFT within an arbitrary time period can easily obtain a customized benchmark with our provided open-source scripts.
4 KEDAS: A Proposal for Real-time Knowledge Editing Alignment
Figure 1: Illustration of our proposed KEDAS framework.
As will be shown in the experiments, previous methods show certain limitations either in edit success and portability or locality on CRAFT. To improve edit success and generalization while keeping a good locality in real-time knowledge editing, we propose KEDAS, an all-round approach consisting of several complementary techniques.
The overall framework of KEDAS is illustrated in Figure 1, including alignment, editing and inference. In alignment, the LLM is aligned with knowledge editing using LoRA. In editing, each edit is converted into diverse forms and stored to the memory module. In the inference phase, edits are retrieved from the memory based on the user query and then filtered by a trained classifier. If there exists relevant edits, the model with trained LoRA adapters will be called for inference. Otherwise, the query will go through the frozen base model. The order of editing and inference is flexible based on practical needs (“edit as you go"), as shown by the bidirectional arrow. For example, we can apply several edits and employ the LLM for inference immediately. If there come further new edits, we can return to the editing phase to conduct these edits and then resume inference.
4.1 Alignment Based on LoRA
LTE Jiang et al. (2024) suffers from a limited locality because it employs the post-alignment model, whose behavior can be significantly different from the original one, to all incoming queries. To address this, we propose the idea of self-adaptive post-alignment inference. To achieve this elegantly, we align the base model via parameter-efficient finetuning with LoRA Hu et al. (2022), which tunes additional adapters during alignment. This costs low computational resources for both training and inference, making KEDAS efficient. We follow LTE to construct alignment data using the training set of CRAFT and details are in Appendix A.
4.2 Diverse Edit Augmentation
Form Content
QA What are Microsoft’s total assets? \$619B.
Declarative Microsoft’s total assets are \$619B.
Aliased MSFT’s total assets are \$619B.
Table 4: A demonstration of diverse edit expressions.
In previous work, each piece of edited knowledge is directly applied to the system as it is, limiting generalization. Since each knowledge can be expressed in various ways, leveraging just one fixed form is inflexible, especially for challenging portability evaluations like in CRAFT.
In this work, we propose diverse edit augmentation, a novel technique that augments each edit by converting it into multiple expressions. The de facto standard form of edits in knowledge editing is a QA pair, typically describing the relation between entities or attributes of them. Based on this observation, we manually design several ways of form augmentation, including declarative and aliased forms, as demonstrated in Table 4.
Concretely, for an edit $e^{t}=(q^{t}_{e},a^{t}_{e})$ at step $t$ 666 For simplicity, we deem the two edits in each CRAFT instance as two separate time steps., its conventional QA form is:
$e^{t}_{\text{qa}}=q^{t}_{e}\oplus a^{t}_{e},$ (7)
where $\oplus$ denotes concatenation. Then, it is augmented as:
$e^{t}_{\text{dec}}=\mathbf{Declarative}(q^{t}_{e},a^{t}_{e}),$ (8)
$e^{t}_{\text{als}}=\mathbf{Aliased}(q^{t}_{e},a^{t}_{e}).$ (9)
Specifically, declarative is to convert the QA pair into a declaration while aliased is to replace specialized terms in the declaration with their aliases for more flexible retrieval.
In experiments, we conduct these augmentations by prompting gpt-4o-mini 777 https://platform.openai.com, an API-based off-the-shelf LLM. The prompts are in Appendix C.
We store the augmented diverse edits into memory for future retrieval:
$\mathcal{M}^{0}=\emptyset,\quad\mathcal{M}^{t}=\mathcal{M}^{t-1}\cup\{e^{t}_{\text{qa}},e^{t}_{\text{dec}},e^{t}_{\text{als}}\},$ (10)
where $\mathcal{M}^{t}$ denotes the memory at time step $t$.
Note that our augmentation methods are specifically designed for the CRAFT dataset. There can be more diverse augmentations like paraphrasing in practice for broader use.
4.3 Filter-based Smart Retriever
During inference, the common practice regarding retrieval in previous knowledge editing work is to directly employ an embedder and perform single-stage retrieval to obtain the top-$k$ hits. This suffers from a poor locality due to the lack of a filter measuring the relevance of queries. In this work, we propose a filter-based smart retriever that pairs the high-recall retrieval with a high-precision filter by precisely identifying relevant queries. In this way, on challenging benchmarks like CRAFT, the post-edit system can keep consistent behaviors for neighboring locality queries while preserving its answers to general queries.
Firstly, we retrieve top-$n$ candidates from the memory using a normal embedding-based retriever. Thanks to our diverse edit augmentation, the memory contains various forms of edits, improving the recall rate. At the time step $t$, for an input query $q$, a set of top-$n$ candidates $\mathcal{C}^{\text{top-}n}$ is retrieved as:
$\mathcal{C}^{\text{top-}n}=\mathbf{Retriever}^{\text{top-}n}(q,\mathcal{M}^{t}).$ (11)
Then, a binary classifier $\theta$ is exploited to predict the relevance between candidates and a query:
$\mathcal{C}^{\text{filtered}}=\{e\mid\mathbf{Filter}_{\theta}(e)=1,e\in\mathcal{C}^{\text{top-}n}\}.$ (12)
Then we only adopt the top-$k$ edits with the highest similarity to the query if the filtered candidate set contains too many edits:
$\mathcal{E}^{*}=\operatorname{arg\,top}_{k}(\mathbf{Similarity}(e,q))\ \text{for }e\in\mathcal{C}^{\text{filtered}}.$ (13)
The filter is trained based on the training set of CRAFT. Details are specified in Appendix D.
4.4 Self-adaptive Post-alignment Inference
In the inference phase, we employ both the original pre-alignment model and the post-alignment adapter, which does not consume extra memory while enabling two inference paths. Each time a query $q$ is requested, the filter-based smart retriever is utilized to recall possible relevant edits in the memory. If there exists a set of relevant edits $\mathcal{E}^{*}$, indicating that the query involves edited knowledge, we fill both $\mathcal{E}^{*}$ and $q$ into the knowledge editing prompt (see Section 2.2) and then feed the prompt into the model with the post-alignment adapters activated. Otherwise, the query directly goes through the pre-alignment model without any processing. The inference strategy is formularized below:
$a=\begin{cases}f_{\Phi}(q)&\text{irrelevant}\\ f_{\Phi+\Delta\Phi(\Theta)}(\mathbf{KEPrompt}(\mathcal{E}^{*},q))&\text{relevant}\\ \end{cases},$ (14)
where $a$ denotes the final answer, $f_{\Phi}$ denotes the original model parametrized by $\Phi$, $f_{\Phi+\Delta\Phi(\Theta)}$ denotes the model with post-alignment adapters, and $\mathbf{KEPrompt}(\cdot,\cdot)$ denotes the knowledge editing prompt template.
Method CRAFT-Statistics CRAFT-Finance
ES $_{\text{temporal}} L$$_{\text{common}}$ $_{\text{reasoning}} P$$_{\text{alias}}$ ES $_{\text{temporal}} L$$_{\text{common}}$ $_{\text{reasoning}} P$$_{\text{alias}}$
LoRA 47.93 0.71 0.00 41.56 47.62 44.19 13.97 0.00 9.81 7.43
ROME 21.59 0.00 12.20 13.67 21.59 4.25 4.00 14.00 7.62 4.40
IKE 99.82 5.49 11.20 43.03 55.73 99.98 27.50 11.60 40.71 57.86
EREN 22.12 91.68 63.20 24.99 22.29 24.37 70.93 65.80 25.48 22.27
WISE 47.54 29.74 81.00 41.26 47.30 32.66 49.62 79.80 34.41 32.48
LTE 97.90 5.44 12.60 68.58 54.14 80.68 27.94 8.40 50.21 60.98
KEDAS (ours) 99.80 100.00 100.00 81.20 62.17 99.85 62.50 100.00 55.95 70.61
Table 5: Main results on CRAFT. The highest and second-highest scores are bolded and underlined, respectively.
5 Experiments
We evaluate representative knowledge editing methods and our proposed KEDAS on CRAFT. Our experiments are based on the framework of EasyEdit Wang et al. (2024b).
5.1 Experimental Setup
Language Model
In our experiments, we focus on Llama-3.1-8B-Instruct Llama Team (2024), a widely used open-source language model.
Evaluated Methods
We evaluate the following knowledge editing methods:
  • LoRA Zhang et al. (2024) directly updates knowledge by LoRA-based finetuning.
  • ROME Meng et al. (2022) locates multilayer perceptron (MLP) weights of related knowledge and writes in new key-value pairs.
  • IKE Zheng et al. (2023) leverages in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting. Due to limited resources, we only evaluate the 8-shot setting.
  • EREN Chen et al. (2024b) prompts LLMs to decide the relevance of retrieved edits and then queries LLMs with or without edits.
  • WISE Wang et al. (2024a) stores edits in a parametric side memory module and utilized a router to decide which memory to go through.
  • LTE Jiang et al. (2024) aligns LLMs with the knowledge editing task by post-training. We adopt paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019) as the retriever and set the number of adopted edits $k$ as 3, the default setting of LTE.
  • KEDAS is our proposed method. We use the same retriever as LTE and adopt bert-base-chinese Devlin et al. (2019) as the binary filter. We set the number of retrieved candidates $n$ as 8 and that of adopted edits $k$ as 3, in line with LTE for fair comparison.
Evaluation Metrics
[…]
5.2 Main Results
The main results on CRAFT are presented in Table 5. Previous methods of knowledge editing show limitations in certain aspects. LoRA and ROME, as traditional methods primarily designed for single editing or limited number of edits, show no advantage when applied to the sequential editing setting on CRAFT. IKE, the retrieval-based approach, suffers from poor locality and mediocre portability, albeit achieves the highest edit success scores. The prompting-based EREN requires specific prompt optimization for different LLMs and can scarcely perform editing with its official default prompt, as can be seen from the low edit success. WISE is parameter-based and suffers from downgrading performance under the sequential setting, with weak performance in most metrics. LTE, with an alignment stage, achieves relatively robust performance in edit success and portability but struggles to achieve locality due to its fixed inference path. These results indicate that CRAFT is a challenging benchmark for existing knowledge editing methods, making it hard for them to achieve a consistently balanced performance.
[…]
In Appendix G, we also compare KEDAS with previous methods on traditional knowledge editing benchmarks. The results show that KEDAS significantly outperform previous methods with balanced performance and confirm the effectiveness of KEDAS even on traditional benchmarks.
Method CRAFT-Statistics CRAFT-Finance
ES $_{\text{temporal}} L$$_{\text{common}}$ $_{\text{reasoning}} P$$_{\text{alias}}$ ES $_{\text{temporal}} L$$_{\text{common}}$ $_{\text{reasoning}} P$$_{\text{alias}}$
KEDAS 99.80 100.00 100.00 81.20 62.17 99.85 62.50 100.00 55.95 70.61
w/o DEA 96.51 100.00 100.00 71.02 59.89 98.10 62.50 100.00 53.90 63.35
w/o FLT 99.80 5.44 11.40 81.20 62.17 100.00 27.48 8.40 55.95 70.61
w/o SPI 99.80 5.68 10.80 81.20 62.17 99.85 27.42 8.40 55.95 70.61
Table 6: Ablation results. The highest scores are bolded.
5.3 Ablation Study
In this section, we assess the indispensability of components of KEDAS, including diverse edit augmentation (DEA), filter (FLT) and self-adaptive post-alignment inference (SPI). The results are presented in Table 6. The exclusion of diverse edit augmentation leads to a decline in edit success and portability, indicating that it contributes to the recall of the retriever. Meanwhile, both the filter and the paradigm of self-adaptive post-alignment inference show great contribution to locality, demonstrating that they are playing important roles in KEDAS. These results confirm that all the core components of KEDAS make contributions to the final performance and are indispensable.
6 Related Work
6.1 Knowledge Editing Benchmarks
Conventionally, knowledge editing benchmarks are created based on temporal changes or counter-fact modifications of world knowledge either from existing QA datasets Levy et al. (2017); Yao et al. (2023) or Wikipedia Cohen et al. (2024a); Zhong et al. (2023), in which Zhong et al. (2023) first propose multi-hop composite portability that involves multiple edits. However, these are all static and suffer from data contamination as time goes by.
Thede et al. (2025) propose WikiBigEdit, a dynamic benchmark based on periodical changes in Wikipedia. However, only a small fraction (3.6%) of its samples support reasoning portability evaluation, limiting its comprehensiveness. Similarly, Tang et al. (2025) propose EvoWiki, a Wiki-based real-time dataset (primarily designed for the RAG), where each instance is associated with reasoning portability tests. Although EvoWiki claims to automatically update with new edits, its implementation and data pipeline remain unpublished.
Besides conventional knowledge editing benchmarks, there have also been some benchmarks designed for special use. Mitchell et al. (2022b) propose ConvSent to edit LLMs’ sentiments on given topics. Ishibashi and Shimodaira (2024) propose Sanitation to forget specific information stored in LLMs and thus address privacy concerns. Huang et al. (2025) propose HalluEditBench to evaluate whether knowledge editing correct hallucinations.
Our proposed CRAFT benchmark leverages statistical and financial data, which are both real-time and inherently dynamic. The strong interdependence among indicators allows each pair of edits to correspond to a meaningful reasoning portability test. This enables comprehensive evaluation across five dimensions: benchmark type, real-time property, real-world grounding, reasoning portability, and composite reasoning portability.
6.2 Knowledge Editing Methods
Parameter-based Editing
This line of work aims to edit model parameters or add extra parameters to perform knowledge editing. Meng et al. (2022)Meng et al. (2023) and Fang et al. (2025) adopt a locate-then-edit manner and aim to precisely edit MLP weights. Mitchell et al. (2022a) adopts meta-learning Hospedales et al. (2022) to predict the change of model parameters. Dong et al. (2022)Huang et al. (2023) and Wang et al. (2024a) add extra parameters to store edits in a parametric form. Mitchell et al. (2022b) trains a separate counter-fact model and does not involve retrieval (thus categorized as parameter-based), failing to fully exploit the capability of LLMs themselves.
Retrieval-based Editing
This line of work stores edits in a memory module and retrieve from it when needed. Zheng et al. (2023) leverages in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting. Chen et al. (2024b) first prompts LLMs to decide the relevance of retrieved edits and then queries LLMs with or without in-context edits. Chen et al. (2024a) converts edits into soft prompt tokens with trained models and then concatenates user prompts to modify the behavior of frozen LLMs. Jiang et al. (2024) aligns LLMs with knowledge editing by post-training with in-context edits. Our proposed KEDAS can also be categorized as retrieval-based.
7 Conclusion
In this work, we introduce CRAFT, a novel automatically updatable real-time benchmark for knowledge editing that focuses on national statistics and finance of China. We evaluate representative knowledge editing methods on CRAFT and reveal that existing methods have certain limitations. We then propose a simple yet effective approach, KEDAS, to better align LLMs with real-time knowledge editing. Experiments validate the effectiveness of KEDAS on CRAFT in all aspects including edit success, locality and portability.
Limitations
First, the CRAFT benchmark focuses on the Chinese language only (but our methodology is universal and can be applied to any other languages) and is limited to the domains of national statistics and finance (but it is easily expandable to domains that data are periodically updated and automatically accessible). Second, the composite reasoning portability of CRAFT is merely 2-hop (but can be expanded to multiple hops via combination of time and indicators). Third, the Finance subset of CRAFT can be updated annually only (but can be improved to be quarterly updated by substituting the annual financial statements to quarter ones). Fourth, only one Llama-3 LLM is employed due to limited time and a limited set of existing knowledge editing methods is evaluated in our experiment due to limited resources and compatibility issues.
Ethical Considerations
Computational Budget
All our experiments are conducted on a machine with Ubuntu 20.04.6 LTS, Intel® Xeon® Silver 4310 CPU and 256G memory. We use one NVIDIA A40 48G GPU for all the experiments. The training of KEDAS takes about 2 hours. The editing of KEDAS takes about 50 minutes on both subsets of CRAFT.
Reproducibility
All the experiments are fully reproducible since all the methods are deterministic and sampling is disabled during LLM generation.
Potential Risks
To the best of our knowledge, there are no potential risks concerning our work.
Scientific Artifacts
Artifact License
[…]
cn-stats MIT
AKShare MIT
BERT Apache
paraphrase-multilingual-MiniLM-L12-v2 Apache
EasyEdit MIT
Llama-3 Llama 3 Community License Agreement
Table 7: Licenses of scientific artifacts we use.
We cite all the creators of scientific artifacts we use in this paper. Licenses of these scientific artifacts are shown in Table 7. Our use of these artifacts is consistent with their intended use.
Privacy and Offense Concerns
Our data source ensures that our CRAFT dataset contains no information that names or uniquely identifies individual people or offensive content.
Appendix A Details of Alignment Data for KEDAS
For each instance of CRAFT’s training set $\mathcal{E}^{t}$, we construct the edit candidate set $\mathcal{E}^{*}$ by keeping the two edits $e_{1}$ and $e_{2}$ of the instance and adding 0-3 (randomly decided) extra retrieved edits with the same retriever as LTE that are different from the two to promote model’s robustness to incorrect retrieved results.
To promote edit success, for each edit query $q_{e}$ and answer $a_{e}$ of either $e_{1}$ or $e_{2}$, the input is $\mathbf{KEPrompt}(\mathcal{E}^{*},q_{e})$ and the target is $a_{e}$.
To promote portability, for each portability query $q_{p}$ and answer $a_{p}$ in $\mathcal{E}^{t}_{p}$, the input is $\mathbf{KEPrompt}(\mathcal{E}^{*},q_{p})$ and the target is $a_{p}$.
To promote locality, for each locality query $q_{l}$ and answer $a_{l}$ in $\mathcal{E}^{t}_{l}$, the in-scope input is $\mathbf{KEPrompt}(\mathcal{E}^{*},q_{l})$ with the target $a_{l}$. We also include an out-of-scope input $q_{l}$ with the target $a_{l}$ that is without any retrieved edits.
In this way, the LLM learns to leverage retrieved edits for relevant queries while keeping the answer unchanged for locality queries.
Appendix B Details of LLM Alignment
Hyperparameter Value
Batch size 1
Gradient accumulation steps 8
Learning rate 1e-4
Epoches 1
Max length 2048
Optimizer AdamW
Scheduler cosine
Warmup ratio 0.1
LoRA rank 8
LoRA alpha 16
LoRA dropout 0
Table 8: Hyper-parameters for LLM finetuning.
We adopt LLaMA-Factory Zheng et al. (2024) to finetune the LLM. The alignment is done on an NVIDIA A40 48GB GPU and takes 2 hours.
The main training hyperparameters are presented in Table 8, most remain unchanged as the default values. Note that we train the LLM for only one epoch because finetuning for multiple epoches will cause over-fitting.
Appendix C Details of Diverse Edit Augmentation
Figure 2: Prompt template for the declarative form.
Figure 3: Prompt template for the aliased form.
The prompt templates of diverse edit augmentation for gpt-4o-mini are presented in Figure 2 and 3. We include manually written in-context examples in the templates to better instruct the model.
Appendix D Details of Training Data for the Filter
For each instance of CRAFT’s training set $\mathcal{E}^{t}$ with $\mathcal{E}^{t}_{e}$, $\mathcal{E}^{t}_{p}$ and $\mathcal{E}^{t}_{l}$ as the set of edits, portability QAs and locality QAs respectively, we simply construct the training data for the filter based on relevance. For any two edits $e=(q_{e},a_{e}),e^{\prime}=(q_{e^{\prime}},a_{e^{\prime}})$ (may be identical) in $\mathcal{E}^{t}_{e}$, the input is $q_{e^{\prime}}\texttt{[sep]}e$ and the target is $1$ (relevant). For any edit $e\in\mathcal{E}^{t}_{e}$ and any portability QA pair $(q_{p},a_{p})\in\mathcal{E}^{t}_{p}$, the input is $q_{p}\texttt{[sep]}e$ and the target is $1$ (relevant). For any edit $e\in\mathcal{E}^{t}_{e}$ and any locality QA pair $(q_{l},a_{l})\in\mathcal{E}^{t}_{l}$, the input is $q_{l}\texttt{[sep]}e$ and the target is 0 (irrelevant). In this way, the filter learns to decide the relevance between the query and the edit.
Appendix E Example from CRAFT-Statistics
  • Prompt 1: 2024年7月中国的货币和准货币(M2)供应量期末值(亿元)是多少? (English translation: What is the period-end value (100 million yuan) of China’s Money and Quasi-Money (M2) supply in July 2024?) Target New: 3033060.78
  • Prompt 2: 2023年7月中国的货币和准货币(M2)供应量期末值(亿元)是多少? (English translation: What is the period-end value (100 million yuan) of China’s Money and Quasi-Money (M2) supply in July 2023?) Target New: 2854031.56
Portability
  • Alias:
    • Prompt: 2024年7月中国的Money and Quasi-Money (M2) Supply, period-end(100 million yuan)是多少? (English translation: What is the Money and Quasi-Money (M2) Supply, period-end (100 million yuan), in July 2024 in China?) Answer: 3033060.78
    • Prompt: 2023年7月中国的Money and Quasi-Money (M2) Supply, period-end(100 million yuan)是多少? (English translation: What is the Money and Quasi-Money (M2) Supply, period-end (100 million yuan), in July 2023 in China?) Answer: 2854031.56
  • Composite Reasoning: Prompt: 2024年7月中国的货币和准货币(M2)供应量期末值(亿元)比2023年7月的高多少? (English translation: How much higher was China’s M2 supply (100 million yuan) in July 2024 than in July 2023?) Answer: 179029.22
Locality
  • Temporal:
    • Prompt: 2022年7月中国的货币和准货币(M2)供应量期末值(亿元)是? (English translation: What was the M2 supply (100 million yuan) at the end of July 2022 in China?)
      Answer: 2578078.57
    • Prompt: 2025年7月中国的货币和准货币(M2)供应量期末值(亿元)是? (English translation: What is the M2 supply (100 million yuan) at the end of July 2025 in China?) Answer: 3299429.06
  • Common-sense:
    • Prompt: 请根据以下材料回答问题。 (略)问题:动物的器官感觉与人的相比有什么不同? (English translation: Read the following passage and answer the question. Question: How do animals’ sensory organs differ from those of humans?) Answer: D
    • Prompt: 请根据以下材料回答问题。 (略)问题:录音中提到能预报风暴的动物是什么? (English translation: Read the following passage and answer the question. Question: Which animal mentioned in the passage can predict storms?) Answer: C
Appendix F Example from CRAFT-Finance
  • Prompt 1: 2024年三友医疗的总资产(亿元)是多少? (English translation: What is the total assets (100 million yuan) of Sanyou Medical in 2024?) Target New: 23.07
  • Prompt 2: 2024年三友医疗的总负债(亿元)是多少? (English translation: What is the total liabilities (100 million yuan) of Sanyou Medical in 2024?) Target New: 2.58
Portability
  • Alias:
    • Prompt: 2024年三友医疗的Total Assets(100 million yuan)是多少? (English translation: What is the total assets (100 million yuan) of Sanyou Medical in 2024?) Answer: 23.07
    • Prompt: 2024年三友医疗的Total Liabilities(100 million yuan)是多少? (English translation: What is the total liabilities (100 million yuan) of Sanyou Medical in 2024?) Answer: 2.58
  • Composite Reasoning: Prompt: 2024年三友医疗的资产负债率(%)是多少? (English translation: What is the debt-to-asset ratio (%) of Sanyou Medical in 2024?)
    Answer: 11.2
Locality
  • Temporal:
    • Prompt: 2023年三友医疗的总资产(亿元)是多少? (English translation: What was the total assets (100 million yuan) of Sanyou Medical in 2023?) Answer: 22.61
    • Prompt: 2023年三友医疗的总负债(亿元)是多少? (English translation: What was the total liabilities (100 million yuan) of Sanyou Medical in 2023?) Answer: 2.19
  • Common-sense:
    • Prompt: 请根据以下材料回答问题。 (略)问题:小羊的结局是什么? (English translation: Read the following passage and answer the question. Question: What happened to the little lamb?) Answer: C
    • Prompt: 请根据以下材料回答问题。 (略)问题:这个故事告诉我们什么道理? (English translation: Read the following passage and answer the question. Question: What moral lesson does this story convey?) Answer: A
Appendix G Results on Traditional Knowledge Editing Benchmarks
Method ZsRE WikiBio $_{\text{recent}} WikiData$$_{\text{counteract}}$
ES L P HM ES L HM ES L P HM ES L P HM
IKE 98.5 43.5 64.4 61.7 96.9 37.9 54.5 97.4 48.8 67.2 65.7 91.3 53.7 61.4 65.4
EREN 32.3 70.7 44.9 44.5 59.4 64.3 61.8 44.4 73.1 37.5 47.7 23.5 75.4 22.9 30.2
WISE 60.0 51.3 40.4 49.3 84.5 95.9 89.8 68.8 94.5 43.1 62.1 47.1 45.4 35.0 41.8
LTE 99.6 55.5 66.7 69.7 97.7 59.7 74.1 99.8 52.9 74.7 70.9 98.3 59.3 73.4 73.8
KEDAS 99.6 90.5 71.4 85.5 97.7 100.0 98.9 99.8 73.9 76.1 81.7 98.3 69.2 77.4 79.9
Table 9: Results on traditional knowledge editing benchmarks. Metrics include edit success (ES), locality (L), portability (P) and harmonic mean (HM) of the previous metrics. The highest scores are bolded.
[…]
The results are presented in Table 9. KEDAS secures the highest harmonic mean scores of edit success, locality and portability on all the four datasets, outperforming previous methods significantly. This further confirms that besides CRAFT, KEDAS is also effective on traditional knowledge editing benchmarks.

### v3

@@
 Abstract
-Knowledge editing aims to modify outdated knowledge in large language models (LLMs) efficiently while retaining their original capabilities.
-Mainstream benchmarks for knowledge editing are predominantly static and fail to keep in pace with the evolving real-world knowledge.
-In this work, we introduce CRAFT, an ever-evolving real-world benchmark for knowledge editing.
-It features well-designed paired edits for composite reasoning, and evaluates models on alias portability as well as temporal and common-sense locality, making it a challenging knowledge editing benchmark on which previous knowledge editing methods hardly achieve balanced performance.
-Towards flexible real-time editing, we propose KEDAS, a novel paradigm of knowledge editing alignment featuring diverse edit augmentation and self-adaptive post-alignment inference, which exhibits significant performance gain on CRAFT compared to previous methods.
-All of our code and data are available at https://anonymous.4open.science/r/CRAFT-KEDAS.
+Knowledge editing aims to modify outdated knowledge in language models efficiently while retaining their original capabilities.
+Mainstream datasets for knowledge editing are predominantly static and fail to keep in pace with the evolving real-world knowledge.
+In this work, we introduce CRAFT, an ever-evolving real-world dataset for knowledge editing.
+It evaluates models on temporal locality, common-sense locality, composite portability and alias portability, providing a comprehensive and challenging evaluation for knowledge editing, on which previous methods hardly achieve balanced performance.
+Towards flexible real-time knowledge editing, we propose KEDAS, a novel paradigm of knowledge editing alignment featuring diverse edit augmentation and self-adaptive post-alignment inference, exhibiting significant performance gain on both CRAFT and traditional datasets compared to previous methods.
+We hope this work may serve as a catalyst for shifting the focus of knowledge editing from static update to dynamic evolution.
+[^1] All of the code and data of this work are publicly available at https://github.com/JamyDon/CRAFT-KEDAS.
 Aligning Language Models with Real-time Knowledge Editing
-Chenming Tang   Yutong Yang   Kexue Wang   Yunfang Wuthanks: Corresponding author.
-National Key Laboratory for Multimedia Information Processing, Peking University MOE Key Laboratory of Computational Linguistics, Peking University School of Computer Science, Peking University {tangchenming, 2300013219}@stu.pku.edu.cn {yytpku, wuyf}@pku.edu.cn
-$\dagger$$\dagger$}\$footnotetext: These authors contributed equally.
+Chenming Tang   Yutong Yang   Kexue Wang   Yunfang Wu National Key Laboratory for Multimedia Information Processing, Peking University School of Computer Science, Peking University {tangchenming, 2300013219}@stu.pku.edu.cn  {yytpku, wuyf}@pku.edu.cn Equal contribution  Corresponding author
 1 Introduction
-Large language models (LLMs) Llama Team (2024); Gemini Team (2025) have been the core of modern natural language processing (NLP).
-Once pre-trained, knowledge is injected into LLMs and becomes their static internal capability Petroni et al.
+Once language models (LMs) are pre-trained, knowledge is injected into them and becomes their static internal capability Petroni et al.
 (2019).
-As time goes by, some knowledge in LLMs inevitably becomes incorrect or out of date, and it is vital to update the outdated knowledge.
-However, retraining a model from scratch is highly costly, especially for downstream users.
+As time goes by, some knowledge in LMs inevitably becomes incorrect or out of date, and it is vital to update the outdated knowledge.
+However, re-training a model frequently is highly costly and usually infeasible.
 To this end, the techniques of knowledge editing have been developed Zhang et al.
-(2024), aimed at editing specific knowledge in LLMs efficiently without heavy re-training.
-There have been a large number of benchmarks to evaluate the effectiveness of knowledge editing methods.
-Nevertheless, most traditional knowledge editing benchmarks like ZsRE Levy et al.
+(2024) to edit specific knowledge in LMs efficiently without heavy re-training.
+There have been a large number of datasets to evaluate the effectiveness of knowledge editing approaches.
+However, most traditional ones like ZsRE Levy et al.
 (2017), MQuAKE Zhong et al.
@@
 Once the datasets are constructed and released, they remain fixed and can not be updated anymore, which is far from real-world knowledge editing applications.
-Recently, some work is devoted to building real-time datasets.
+Recently, there have been a few real-time datasets released.
 For example, WikiBigEdit Thede et al.
-(2025) collects real-world changes in Wikipedia, but it requires processing massive Wiki data (usually hundreds of gigabytes) from the Internet, which is costly and inconvenient in practice.
-Meanwhile, the automatically collected dataset lacks necessary filtering and thus suffers from severe sparsity (for example, only 3.6% of the collected data have an entry of portability evaluation).
+(2025) collects real-world changes in Wikipedia, but it requires processing massive Wiki data (hundreds of gigabytes) from the Internet, which is costly and inconvenient in practice.
+Meanwhile, the automatically collected dataset lacks necessary filtering and thus suffers from severe sparsity (for instance, only 3.6% of the collected data have an entry of portability evaluation).
 EvoWiki Tang et al.
-(2025) is another real-time dataset also based on Wikipedia, but it focuses on retrieval-augmented generation (RAG) and continual learning.
-Overall, there still lacks a high-quality, easily-updated benchmark for real-world and real-time knowledge editing.
-Benchmark Real- time Real- world Reasoning Portability Composite Reasoning Portability
+(2025) is another real-time dataset based on Wikipedia, but it focuses on retrieval-augmented generation (RAG) and continual learning and does not support comprehensive evaluation for knowledge editing due to lack of crucial aspects like locality.
+Overall, there still lacks a high-quality, easily-updated dataset for real-world and real-time knowledge editing.
+Dataset RT RW RP CRP
 ZsRE
@@
 CRAFT (Ours)
-Table 1: Comparison of knowledge editing benchmarks.
-Composite reasoning portability evaluates a model’s ability to integrate multiple distinct edits into a single reasoning query, whereas reasoning portability tests reasoning chains derived from a single edit.
-In this work, we introduce CRAFT (Chinese Real-time statistics And Finance knowledge ediTing benchmark), an ever-evolving real-world benchmark for knowledge editing in Chinese.
+Table 1: Comparison of knowledge editing datasets.
+RT refers to “real-time" and RW denotes “real-world".
+RP (reasoning portability) tests reasoning chains derived from a single edit, whereas CRP (composite reasoning portability) evaluates a model’s ability to integrate multiple distinct edits into a single reasoning query.
+In this work, we introduce CRAFT (Chinese Real-time statistics And Finance knowledge editing dataseT), an ever-evolving real-world dataset for knowledge editing in Chinese.
 It leverages publicly available official data that are continuously updated, ensuring both temporal freshness and real-world applications.
-Furthermore, it is organized in well-designed paired edits, each serving as a composite reasoning test to evaluate models’ ability to integrate multiple related factual updates.
-Moreover, CRAFT supports evaluations on alias portability and temporal and common-sense locality, providing a comprehensive assessment of model adaptability and robustness under dynamic knowledge changes.
+CRAFT is organized in well-designed paired edits, each serving as a composite reasoning test to evaluate models’ ability to integrate multiple related factual updates.
+Meanwhile, it supports evaluations on alias portability and temporal and common-sense locality, providing a comprehensive assessment of LMs’ adaptability and robustness under dynamic knowledge changes.
 The differences between CRAFT and previous datasets are summarized in Table 1.
-We assess the exposure rate of CRAFT and existing benchmarks with five different LLMs, either open-source or closed-source.
-The results demonstrate that a large portion of knowledge in existing benchmarks has been known to LLMs while the exposure rate of CRAFT is nearly zero on all the five models.
-This reveals the disadvantages of static datasets and validates the real-time property of CRAFT, which makes it challenging to LLMs.
-We evaluate a suite of representative knowledge editing methods on CRAFT and demonstrate that current approaches exhibit inherent limitations.
-For example, parameter-based approaches like ROME Meng et al.
+We perform exposure analysis on CRAFT and other evaluation datasets with five different LMs.
+The results indicate that a large portion of knowledge in existing datasets has been known to LMs while the exposure rate of CRAFT is nearly zero.
+This validates the real-time property of CRAFT, which makes it challenging and fair.
+We evaluate a suite of representative knowledge editing methods on CRAFT and demonstrate current approaches exhibit inherent limitations.
+Parameter-based approaches like ROME Meng et al.
 (2022) and WISE Wang et al.
-(2024a) suffer from gradual model degradation and struggle to achieve successful editing in the setting of sequential editing.
+(2024a) suffer from gradual model degradation and scarcely edit successfully when new knowledge comes sequentially and continually.
 Simple retrieval-based approaches like IKE Zheng et al.
@@
 LTE Jiang et al.
-(2024), which aligns LLMs with knowledge editing via post-training, shows weak locality due to overfitting on irrelevant queries.
+(2024), which aligns LMs with knowledge editing via finetuning, shows weak locality due to overfitting on irrelevant queries.
 […]
 Our contributions are as follows:
-  • We introduce CRAFT, a high-quality, easily-updated and fully open-source benchmark for real-time knowledge editing.
-  • We evaluate existing knowledge editing methods on CRAFT and reveal the limitations of these methods in real-time knowledge editing.
-  • We propose KEDAS, a novel paradigm of aligning LLMs with real-time knowledge editing featuring diverse edit augmentation and self-adaptive post-alignment inference, enabling dynamic routing between pre- and post-alignment LLMs, which significantly outperforms existing methods on CRAFT.
+  • We reveal the leakage issue of existing datasets and introduce CRAFT, an open-source, contamination-free and easily-updated dataset for real-time knowledge editing.
+  • We evaluate existing knowledge editing methods on CRAFT and reveal the limitations of these methods in real-time knowledge editing, especially the tradeoff among edit success, locality and portability.
+  • We propose KEDAS, a novel paradigm of aligning LMs with real-time knowledge editing which significantly outperforms existing methods with balanced performance.
+This paper is organized as follows.
+We introduce the preliminary of knowledge editing in § 2.
+Then, we present our CRAFT dataset in § 3 and propose KEDAS for real-time knowledge editing in § 4.
+After that, we conduct empirical experiments in § 5.
+Finally, we discuss related work in § 6 and conclude our study in § 7.
 2 Preliminary
 2.1 Task Formulation
-An LLM system $f$ can be regarded as a function $f:\mathcal{Q}\mapsto\mathcal{A}$ mapping an query $q\in\mathcal{Q}$ to its output answer $a=f(q)\in\mathcal{A}$.
-Knowledge editing aims to change the behavior of the LLM after modifying some knowledge. Edits of knowledge are usually in the form of query-answer (QA) pairs:
+An LM system $f$ can be regarded as a function $f:\mathcal{Q}\mapsto\mathcal{A}$ mapping an query $q\in\mathcal{Q}$ to its output answer $a=f(q)\in\mathcal{A}$.
+Knowledge editing aims to change the behavior of the LM after modifying some knowledge. Edits of knowledge are usually in the form of query-answer (QA) pairs:
 $\mathcal{E}=\{e^{t}\}^{N}_{t=1}=\{(q^{t}_{e},a^{t}_{e})\}^{N}_{t=1},$ (1)
 where $q^{t}_{e}$ is an input query triggering knowledge (e.g., The current US president is), $a^{t}_{e}$ is the corresponding target of editing (e.g., Donald Trump), $t$ is the index of each edit, and $N$ denotes the total number of edits.
-To assess the efficacy of editing, the post-edit LLM $f^{*}$ is evaluated via the following criteria Zhang et al. (2024):
+To assess the efficacy of knowledge editing, the post-edit LM $f^{*}$ is evaluated via the following criteria Zhang et al. (2024):
 Edit success
@@
 $f^{t}=\mathbf{Edit}(f,q^{t}_{e},a^{t}_{e}),\quad 1\leq t\leq N.$ (5)
-This traditional setting is widely employed by earlier work. It is far from real-world applications, since only one edit can be applied each time.
+This traditional setting is widely employed by earlier work. It is far from real-world applications, since only one edit can be applied.
 Sequential Editing
-In sequential editing, edits are applied step by step (let $f^{0}$ denote $f$ for consistency):
+In sequential editing, edits are applied step by step in a continual manner ($f^{0}$ denotes $f$ for consistency):
 $f^{t}=\mathbf{Edit}(f^{t-1},q^{t}_{e},a^{t}_{e}),\quad 1\leq t\leq N.$ (6)
@@
 We focus on this setting in this work because it is closer to practical applications, especially the real-time knowledge editing setting.
-Knowledge Editing Alignment
+2.3 Knowledge Editing Alignment
 LTE Jiang et al.
-(2024) relates knowledge editing to LLM alignment, aligning LLMs with ever-changing real-time knowledge edits.
-During alignment, LLMs’ capabilities of knowledge updating are enlightened by an editing prompt:
+(2024) relates knowledge editing to LM alignment, aligning LMs with ever-changing real-time knowledge edits.
+During alignment, LMs’ capabilities of knowledge updating are enlightened by an editing prompt:
     [Updated Information] {edit}
@@
 The training data consist of in-scope and out-of-scope queries with or without the specific edit to improve edit efficacy while keeping locality.
-In the inference phase, LLMs are required to conduct on-the-fly and streaming knowledge editing by retrieving relevant edits to the query from the stored memory.
+In the inference phase, LMs are required to conduct on-the-fly and streaming knowledge editing by retrieving relevant edits to the query from the stored memory.
 Dataset GPT-3.5 Mistral-7B Llama-3.1-8B DeepSeek-chat Doubao-1.5-pro-32k Size
@@
 ZsRE Levy et al. (2017) 43.98% 23.48% 22.45% 52.47% 26.77% 6,000
-MQuAKE Zhong et al. (2023) 35.46% 76.21% 46.11% 42.58% 76.09% 5,604
-EvoWiki_evolved Tang et al. (2025) 8.07% 16.22% 6.43% 10.24% 5.08% 1,338
-EvoWiki_stable Tang et al. (2025) 28.45% 42.37% 25.90% 39.76% 12.32% 1,494
-EvoWiki_uncharted Tang et al. (2025) 8.19% 22.62% 7.39% 16.02% 3.43% 1,136
+MQuAKE-T Zhong et al. (2023) 35.46% 76.21% 46.11% 42.58% 76.09% 5,604
+[…]
+[…]
+[…]
 CRAFT-Statistics (Ours) 0.01% 0.00% 0.30% 1.80% 0.99% 4,468
@@
 3 The CRAFT Dataset
-We introduce CRAFT, a real-time, real-world benchmark designed to evaluate knowledge editing under dynamically evolving factual contexts.
-Unlike static benchmarks that rely on outdated or widely exposed Wikipedia facts, CRAFT continuously updates with publicly available official data sources, ensuring fairness and real-world relevance.
-3.1 Dataset Overview
-CRAFT consists of two complementary subsets: CRAFT-Statistics and CRAFT-Finance, representing stable and highly correlated factual domains that evolve periodically.
-Example instances from the two subsets are shown in Appendix E and F respectively.
+We introduce CRAFT, a real-time and real-world dataset designed to evaluate knowledge editing under dynamically evolving factual contexts.
+Unlike static datasets that rely on outdated or widely exposed Wikipedia facts, CRAFT continuously updates with publicly available official data sources, ensuring fairness and real-world relevance.
+3.1 Data Sources
+CRAFT consists of two complementary subsets, namely CRAFT-Statistics and CRAFT-Finance, representing two highly correlated real-world domains that evolve periodically.
+Example instances from the two subsets are shown in Appendix G and H respectively.
 CRAFT-Statistics
-We collect monthly statistical reports from the official National Bureau of Statistics of China 222 https://data.stats.gov.cn/, covering July 2024 to June 2025 in the current version used in this work.
-The dataset spans four major domains—finance, fiscal affairs, telecommunications, and transportation—containing 221 indicators and 4,468 data points.
-Our data collection pipeline, leveraging the cn-stats API 333 https://github.com/songjian/cnstats, supports automatic temporal updates, enabling effortless retrieval of the most recent monthly data.
+We collect monthly statistical reports from the National Bureau of Statistics of China [^3] https://data.stats.gov.cn/, covering July 2024 to June 2025 in the current version used in this work.
+The dataset spans four major domains – finance, fiscal affairs, telecommunications, and transportation – containing 221 indicators and 4,468 data points.
+Our data collection pipeline, leveraging the cn-stats API [^4] https://github.com/songjian/cnstats, supports automatic temporal updates, enabling effortless collection of the most recent monthly data.
 CRAFT-Finance
-CRAFT-Finance includes annual financial statements of 390 publicly listed Chinese companies (2023–2024), sourced from Eastmoney Financial Database 444 https://data.eastmoney.com/ via the AKShare API 555 https://github.com/akfamily/akshare.
+CRAFT-Finance includes annual financial statements of 390 publicly listed Chinese companies (2023–2024), collected via the AKShare API [^5] https://github.com/akfamily/akshare.
 It features financial indicators such as debt ratio, shareholder equity, and operating cash flow, supporting yearly updates to reflect the latest financial disclosures.
 3.2 Evaluation Metrics
-To comprehensively assess model adaptability, besides the conventional edit success (ES) as described in Section 2.1, CRAFT provides four evaluation dimensions:
+To comprehensively assess model adaptability, CRAFT provides four evaluation dimensions besides the conventional edit success (ES) as described in § 2.1:
 […]
@@
 The corresponding composite reasoning query and answer are:
-$q_{p}$ $=\text{2024 Company A's debt ratio}$
-$a_{p}$ $=0.5$
-CRAFT-Finance includes ten types of such combinations, demonstrating the generality and practical relevance of composite reasoning portability.
+$q_{p}$ $:\text{2024 Company A's debt ratio}$
+$a_{p}$ $:0.5$
+CRAFT-Finance includes 10 types of such combinations, demonstrating the generality and practical relevance of composite reasoning portability.
 For CRAFT-Statistics, composite portability is constructed by pairing two edits of the same indicator from adjacent years.
@@
 […]
+3.3 Exposure Analysis
+A perfect evaluation dataset should be unleaked, with a low rate of knowledge exposed to LMs (i.e., exposure rate).
+If a large portion of knowledge in a knowledge editing dataset has already been known by LMs, the evaluation results on it can be inaccurate, unfair, or even misleading.
+We evaluate the exposure rates of CRAFT and existing benchmarks across five representative LMs by directly prompting with the questions in the datasets and counting how many questions can be correctly answered.
+As shown in Table 2, Due to its temporal freshness and domain specificity, CRAFT exhibits extremely low exposure rates (nearly 0%) across all the five LMs, mitigating data contamination and ensuring fair evaluation.
+Meanwhile, for other datasets, a great portion of the knowledge has been known to the LMs, making them less challenging and less fair in practice.
+This reveals the advantage of CRAFT as a truly real-time dataset.
 3.4 Dataset Statistics
-Table 2 compares the exposure ratio of CRAFT and existing benchmarks across five representative LLMs.
-Due to its temporal freshness and domain specificity, CRAFT exhibits extremely low exposure rates across all the five LLMs, mitigating data contamination and ensuring fair evaluation.
-Meanwhile, for other datasets, a great portion of the knowledge has been known to LLMs, making them less challenging and less fair in practice.
-Dataset Edit $\mathbf{P}_{\text{reasoning}}$ $\mathbf{P}_{\text{alias}}$ $\mathbf{L}_{\text{temporal}}$ $\mathbf{L}_{\text{common}}$
+Subset Edit $\mathbf{P}_{\text{reasoning}}$ $\mathbf{P}_{\text{alias}}$ $\mathbf{L}_{\text{temporal}}$ $\mathbf{L}_{\text{common}}$
 CRAFT- Statistics 4,468 2,234 4,468 4,468 4,468
 CRAFT- Finance 7,800 3,900 7,800 7,800 7,800
-Table 3: Number of instances per evaluation type in CRAFT. Composite reasoning uses edit pairs; other metrics are per edit.
-Table 3 shows instance counts in CRAFT sub-datasets.
+Table 3: The numbers of instances in different evaluation types of CRAFT. Composite reasoning uses edit pairs; other metrics are per edit.
+Table 3 shows the numbers of instances in CRAFT’s subsets.
 Composite reasoning portability is evaluated per edit pair, and other metrics per edit.
-For experiments, we split CRAFT into training and test sets.
+For experiments, we split each subset of CRAFT into training and test sets.
 The test set contains 500 instances from the two subsets each (1,000 instances in total) and the training set contains the remaining instances.
-3.4 Data Summary
-CRAFT offers a scalable and continuously evolving benchmark reflecting real-world knowledge updates.
-Its dynamic nature, factual authenticity, and comprehensive reasoning evaluation make it a robust foundation for future research on reliable and interpretable knowledge editing.
-Anyone who wants to construct CRAFT within an arbitrary time period can easily obtain a customized benchmark with our provided open-source scripts.
+3.5 Discussion
+CRAFT is a scalable and continuously evolving dataset reflecting real-world knowledge updates.
+Its dynamic nature, factual authenticity, and comprehensive reasoning evaluation make it a robust foundation for future research on reliable knowledge editing.
+Anyone who wants to construct CRAFT within an arbitrary time period can easily obtain a customized dataset with our provided open-source scripts.
+Through CRAFT, we also demonstrate that researchers can make full use of public available data sources (like public APIs) besides the widely used Wikipedia data to create datasets with real-world and real-time properties and scalability that can help evaluate, train or develop more scalable intelligent systems without much human labor.
 4 KEDAS: A Proposal for Real-time Knowledge Editing Alignment
 Figure 1: Illustration of our proposed KEDAS framework.
-As will be shown in the experiments, previous methods show certain limitations either in edit success and portability or locality on CRAFT.
-To improve edit success and generalization while keeping a good locality in real-time knowledge editing, we propose KEDAS, an all-round approach consisting of several complementary techniques.
+As will be shown in the experiments (§ 5), previous methods show certain limitations either in recall (i.e., edit success and portability) or precision (i.e., locality) on CRAFT.
+To improve edit success and generalization while keeping a good locality in real-time knowledge editing, we propose KEDAS, an all-round approach consisting of several complementary designs.
 The overall framework of KEDAS is illustrated in Figure 1, including alignment, editing and inference.
-In alignment, the LLM is aligned with knowledge editing using LoRA.
+In alignment, the LM is aligned with knowledge editing using LoRA.
 In editing, each edit is converted into diverse forms and stored to the memory module.
 In the inference phase, edits are retrieved from the memory based on the user query and then filtered by a trained classifier.
-If there exists relevant edits, the model with trained LoRA adapters will be called for inference.
+If there exist relevant edits, the model with trained LoRA adapters will be called for inference.
 Otherwise, the query will go through the frozen base model.
-The order of editing and inference is flexible based on practical needs (“edit as you go"), as shown by the bidirectional arrow.
-For example, we can apply several edits and employ the LLM for inference immediately.
-If there come further new edits, we can return to the editing phase to conduct these edits and then resume inference.
+The order of editing and inference is flexible based on practical needs (“edit as you go”), as shown by the bidirectional arrow.
+For example, we may apply several edits and employ the LM for inference immediately.
+If there come further new edits, we can return to the editing phase to perform these edits and then resume inference.
 4.1 Alignment Based on LoRA
@@
 In previous work, each piece of edited knowledge is directly applied to the system as it is, limiting generalization.
-Since each knowledge can be expressed in various ways, leveraging just one fixed form is inflexible, especially for challenging portability evaluations like in CRAFT.
+Since each knowledge can be expressed in various ways, leveraging just one fixed form is inflexible, especially for challenging portability evaluations like that in CRAFT.
 In this work, we propose diverse edit augmentation, a novel technique that augments each edit by converting it into multiple expressions.
@@
 Based on this observation, we manually design several ways of form augmentation, including declarative and aliased forms, as demonstrated in Table 4.
-Concretely, for an edit $e^{t}=(q^{t}_{e},a^{t}_{e})$ at step $t$ 666 For simplicity, we deem the two edits in each CRAFT instance as two separate time steps., its conventional QA form is:
+Concretely, for an edit $e^{t}=(q^{t}_{e},a^{t}_{e})$ at step tt [^6] For simplicity, we deem the two edits in each CRAFT instance as two separate time steps., its conventional QA form is:
 $e^{t}_{\text{qa}}=q^{t}_{e}\oplus a^{t}_{e},$ (7)
@@
 Specifically, declarative is to convert the QA pair into a declaration while aliased is to replace specialized terms in the declaration with their aliases for more flexible retrieval.
-In experiments, we conduct these augmentations by prompting gpt-4o-mini 777 https://platform.openai.com, an API-based off-the-shelf LLM. The prompts are in Appendix C.
+In experiments, we conduct these augmentations by prompting an API-based off-the-shelf LM.
 We store the augmented diverse edits into memory for future retrieval:
@@
 In this work, we propose a filter-based smart retriever that pairs the high-recall retrieval with a high-precision filter by precisely identifying relevant queries.
-In this way, on challenging benchmarks like CRAFT, the post-edit system can keep consistent behaviors for neighboring locality queries while preserving its answers to general queries.
-Firstly, we retrieve top-$n$ candidates from the memory using a normal embedding-based retriever.
+In this way, on challenging datasets like CRAFT, the post-edit system can keep consistent behaviors for neighboring locality queries and preserve its original answers to general queries.
+First, we retrieve top-$n$ candidates from the memory using a normal embedding-based retriever.
 Thanks to our diverse edit augmentation, the memory contains various forms of edits, improving the recall rate.
@@
 Then we only adopt the top-$k$ edits with the highest similarity to the query if the filtered candidate set contains too many edits:
-$\mathcal{E}^{*}=\operatorname{arg\,top}_{k}(\mathbf{Similarity}(e,q))\ \text{for }e\in\mathcal{C}^{\text{filtered}}.$ (13)
-The filter is trained based on the training set of CRAFT. Details are specified in Appendix D.
+$\mathcal{E}^{*}=\operatorname{arg\,top}_{k}(\mathbf{Sim}(e,q))\ \text{for }e\in\mathcal{C}^{\text{filtered}}.$ (13)
 4.4 Self-adaptive Post-alignment Inference
@@
 Each time a query $q$ is requested, the filter-based smart retriever is utilized to recall possible relevant edits in the memory.
-If there exists a set of relevant edits $\mathcal{E}^{*}$, indicating that the query involves edited knowledge, we fill both $\mathcal{E}^{*}$ and $q$ into the knowledge editing prompt (see Section 2.2) and then feed the prompt into the model with the post-alignment adapters activated.
+If there exists a set of relevant edits $\mathcal{E}^{*}$, indicating that the query may involve edited knowledge, we fill both $\mathcal{E}^{*}$ and $q$ into the knowledge editing prompt (see § 2.3) and then pass the prompt to the model with the post-alignment adapters activated.
 Otherwise, the query directly goes through the pre-alignment model without any processing.
@@
 $a=\begin{cases}f_{\Phi}(q)&\text{irrelevant}\\ f_{\Phi+\Delta\Phi(\Theta)}(\mathbf{KEPrompt}(\mathcal{E}^{*},q))&\text{relevant}\\ \end{cases},$ (14)
-where $a$ denotes the final answer, $f_{\Phi}$ denotes the original model parametrized by $\Phi$, $f_{\Phi+\Delta\Phi(\Theta)}$ denotes the model with post-alignment adapters, and $\mathbf{KEPrompt}(\cdot,\cdot)$ denotes the knowledge editing prompt template.
-Method CRAFT-Statistics CRAFT-Finance
+where $a$ denotes the final answer, $f_{\Phi}$ denotes the original model parametrized by $\Phi$, $f_{\Phi+\Delta\Phi(\Theta)}$ denotes the model with post-alignment adapters $\Delta\Phi(\Theta)$, and $\mathbf{KEPrompt}(\cdot,\cdot)$ denotes the knowledge editing prompt template.
+4.5 Discussion
+KEDAS borrows the idea of inference-time retrieval from EREN Chen et al.
+(2024b) and the idea of knowledge editing alignment from LTE Jiang et al.
+(2024).
+The novelty of KEDAS lies in the two unique designs: diverse edit augmentation and filter-based smart retriever paired with self-adaptive post-alignment inference.
+The former follows the idea that edits can go beyond their initial forms, while the latter demonstrates the insight that we do not have to perform knowledge editing for irrelevant queries and a lightweight flexible mechanism leads to more balanced performance.
+Model Method CRAFT-Statistics CRAFT-Finance
 ES $_{\text{temporal}} L$$_{\text{common}}$ $_{\text{reasoning}} P$$_{\text{alias}}$ ES $_{\text{temporal}} L$$_{\text{common}}$ $_{\text{reasoning}} P$$_{\text{alias}}$
-LoRA 47.93 0.71 0.00 41.56 47.62 44.19 13.97 0.00 9.81 7.43
+Llama-3.1 LoRA 47.93 0.71 0.00 41.56 47.62 44.19 13.97 0.00 9.81 7.43
 ROME 21.59 0.00 12.20 13.67 21.59 4.25 4.00 14.00 7.62 4.40
@@
 LTE 97.90 5.44 12.60 68.58 54.14 80.68 27.94 8.40 50.21 60.98
-KEDAS (ours) 99.80 100.00 100.00 81.20 62.17 99.85 62.50 100.00 55.95 70.61
+KEDAS 99.80 100.00 100.00 81.20 62.17 99.85 62.50 100.00 55.95 70.61
+Qwen2.5 LoRA 43.81 18.86 0.00 39.48 35.91 24.73 11.80 0.00 31.59 32.89
+ROME 35.06 32.79 0.00 29.00 26.90 9.75 2.95 0.00 14.39 9.34
+IKE 96.82 19.44 82.40 48.48 42.02 97.96 28.78 81.80 39.27 46.16
+EREN 28.32 61.08 92.60 34.55 30.67 31.25 42.43 93.00 33.69 29.53
+WISE 40.72 43.46 94.40 34.57 38.39 28.07 51.43 95.00 28.67 21.25
+LTE 99.95 8.38 6.60 77.67 58.92 99.96 31.56 5.20 51.86 60.18
+KEDAS 99.95 100.00 100.00 80.88 58.59 99.88 66.47 100.00 51.55 65.97
 Table 5: Main results on CRAFT. The highest and second-highest scores are bolded and underlined, respectively.
 5 Experiments
-We evaluate representative knowledge editing methods and our proposed KEDAS on CRAFT. Our experiments are based on the framework of EasyEdit Wang et al. (2024b).
 5.1 Experimental Setup
-Language Model
-In our experiments, we focus on Llama-3.1-8B-Instruct Llama Team (2024), a widely used open-source language model.
-Evaluated Methods
-We evaluate the following knowledge editing methods:
-  • LoRA Zhang et al. (2024) directly updates knowledge by LoRA-based finetuning.
-  • ROME Meng et al. (2022) locates multilayer perceptron (MLP) weights of related knowledge and writes in new key-value pairs.
-  • IKE Zheng et al. (2023) leverages in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting. Due to limited resources, we only evaluate the 8-shot setting.
-  • EREN Chen et al. (2024b) prompts LLMs to decide the relevance of retrieved edits and then queries LLMs with or without edits.
-  • WISE Wang et al. (2024a) stores edits in a parametric side memory module and utilized a router to decide which memory to go through.
-  • LTE Jiang et al.
-(2024) aligns LLMs with the knowledge editing task by post-training.
-We adopt paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019) as the retriever and set the number of adopted edits $k$ as 3, the default setting of LTE.
-  • KEDAS is our proposed method.
-We use the same retriever as LTE and adopt bert-base-chinese Devlin et al.
-(2019) as the binary filter.
-We set the number of retrieved candidates $n$ as 8 and that of adopted edits $k$ as 3, in line with LTE for fair comparison.
-Evaluation Metrics
+5.1.1 Evaluation
 […]
-5.2 Main Results
-The main results on CRAFT are presented in Table 5.
+[…]
+In the main experiments, we focus on sequential editing for both evaluation datasets.
+5.1.2 Language Models
+For CRAFT, we use Llama-3.1-8B-Instruct Llama Team (2024) and Qwen2.5-3B-Instruct Qwen (2025), both from widely used herds of open-source LMs (Llama-3.1 and Qwen2.5) and with different scales.
+For KnowEdit, we focus on the same Llama-3.1 model and provide the results on two other LMs in Appendix I.
+5.1.3 Evaluated Methods
+We evaluate the following knowledge editing methods: (1) LoRA Zhang et al.
+(2024) directly updates knowledge by LoRA-based finetuning.
+(2) ROME Meng et al.
+(2022) locates multilayer perceptron weights of related knowledge and writes in new key-value pairs.
+(3) IKE Zheng et al.
+(2023) leverages in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting.
+(4) EREN Chen et al.
+(2024b) prompts LMs to decide the relevance of retrieved edits and then queries LMs with or without edits.
+(5) WISE Wang et al.
+(2024a) stores edits in a parametric side memory module and utilized a router to decide which memory to go through.
+(6) LTE Jiang et al.
+(2024) aligns LMs with the knowledge editing task by post-training.
+(7) KEDAS is our proposed method.
+Due to compatibility, cost, and performance issues, we do not include some other baselines as discussed in Appendix F.
+5.1.4 Implementation Details
+Our experiments are based on the framework of EasyEdit Wang et al. (2024b). For most of the baselines, we follow the default hyper-parameter settings provided by EasyEdit.
+For IKE, we only evaluate the 8-shot setting because of limited resources.
+For both LTE and KEDAS, we use paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019) as the retriever and set the number of adopted edits $k$ as 3, the default setting of LTE.
+For KEDAS, we adopt bert-base-chinese Devlin et al.
+(2019) as the binary filter, which is trained based on the training set of CRAFT with details specified in Appendix D.
+We set the number of retrieved candidates $n$ as 8 and that of adopted edits $k$ as 3, in line with LTE.
+The diverse edit augmentation of KEDAS is done by prompting gpt-4o-mini [^7] https://platform.openai.com and the prompts are in Appendix C.
+5.2 Results on CRAFT
+The results on CRAFT are presented in Table 5.
 Previous methods of knowledge editing show limitations in certain aspects.
-LoRA and ROME, as traditional methods primarily designed for single editing or limited number of edits, show no advantage when applied to the sequential editing setting on CRAFT.
-IKE, the retrieval-based approach, suffers from poor locality and mediocre portability, albeit achieves the highest edit success scores.
-The prompting-based EREN requires specific prompt optimization for different LLMs and can scarcely perform editing with its official default prompt, as can be seen from the low edit success.
+LoRA and ROME, as traditional methods primarily designed for single editing or a limited number of edits, show no advantage when applied to the sequential editing setting on CRAFT.
+IKE, the retrieval-based approach, suffers from poor locality and mediocre portability, albeit achieves outstanding edit success.
+The prompting-based EREN requires specific prompt optimization for different LMs and can scarcely perform editing with its default prompt, as can be seen from the low edit success.
 WISE is parameter-based and suffers from downgrading performance under the sequential setting, with weak performance in most metrics.
 LTE, with an alignment stage, achieves relatively robust performance in edit success and portability but struggles to achieve locality due to its fixed inference path.
-These results indicate that CRAFT is a challenging benchmark for existing knowledge editing methods, making it hard for them to achieve a consistently balanced performance.
+These indicate that CRAFT is challenging for existing knowledge editing methods, making it hard to achieve a consistently balanced performance.
 […]
-In Appendix G, we also compare KEDAS with previous methods on traditional knowledge editing benchmarks.
-The results show that KEDAS significantly outperform previous methods with balanced performance and confirm the effectiveness of KEDAS even on traditional benchmarks.
+5.3 Results on KnowEdit
+Method ZsRE WikiBio $_{\text{recent}} WikiData$$_{\text{counteract}}$
+ES L P HM ES L HM ES L P HM ES L P HM
+IKE 98.5 43.5 64.4 61.7 96.9 37.9 54.5 97.4 48.8 67.2 65.7 91.3 53.7 61.4 65.4
+EREN 32.3 70.7 44.9 44.5 59.4 64.3 61.8 44.4 73.1 37.5 47.7 23.5 75.4 22.9 30.2
+WISE 60.0 51.3 40.4 49.3 84.5 95.9 89.8 68.8 94.5 43.1 62.1 47.1 45.4 35.0 41.8
+LTE 99.6 55.5 66.7 69.7 97.7 59.7 74.1 99.8 52.9 74.7 70.9 98.3 59.3 73.4 73.8
+KEDAS 99.6 90.5 71.4 85.5 97.7 100.0 98.9 99.8 73.9 76.1 81.7 98.3 69.2 77.4 79.9
+Table 6: Results on KnowEdit. Metrics include edit success (ES), locality (L), portability (P) and harmonic mean (HM) of the previous metrics. The highest scores are bolded.
+The results on KnowEdit are shown in Table 6.
+KEDAS exhibits outstanding performance in all the metrics and secures the highest harmonic mean (HM) scores across all the datasets.
+On average, its HM outperforms WISE by +26.9 and surpasses LTE by +18.2.
+As an alignment-based approach, KEDAS is superior especially in locality compared with LTE, with an improvement of +31.7.
+On KnowEdit, we also further experiment with other LMs and under the traditional setting of single editing.
+These results can be found in Appendix I, where KEDAS also secures consistent advantage with balanced performance.
+5.4 Ablation Study
 Method CRAFT-Statistics CRAFT-Finance
@@
 w/o SPI 99.80 5.68 10.80 81.20 62.17 99.85 27.42 8.40 55.95 70.61
-Table 6: Ablation results. The highest scores are bolded.
-5.3 Ablation Study
+Table 7: Ablation results on CRAFT. The highest scores are bolded.
 In this section, we assess the indispensability of components of KEDAS, including diverse edit augmentation (DEA), filter (FLT) and self-adaptive post-alignment inference (SPI).
@@
 6 Related Work
-6.1 Knowledge Editing Benchmarks
-Conventionally, knowledge editing benchmarks are created based on temporal changes or counter-fact modifications of world knowledge either from existing QA datasets Levy et al.
+6.1 Related Evaluation Datasets
+Conventionally, knowledge editing datasets are created based on temporal changes or counter-fact modifications of world knowledge either from existing QA datasets Levy et al.
 (2017); Yao et al.
@@
 (2023), in which Zhong et al.
-(2023) first propose multi-hop composite portability that involves multiple edits.
+(2023) first propose multi-hop composite portability.
 However, these are all static and suffer from data contamination as time goes by.
 Thede et al.
-(2025) propose WikiBigEdit, a dynamic benchmark based on periodical changes in Wikipedia.
+(2025) propose WikiBigEdit, a dynamic dataset based on periodical changes in Wikipedia.
 However, only a small fraction (3.6%) of its samples support reasoning portability evaluation, limiting its comprehensiveness.
 Similarly, Tang et al.
-(2025) propose EvoWiki, a Wiki-based real-time dataset (primarily designed for the RAG), where each instance is associated with reasoning portability tests.
+(2025) propose EvoWiki, a Wiki-based real-time dataset primarily designed for RAG.
 Although EvoWiki claims to automatically update with new edits, its implementation and data pipeline remain unpublished.
-Besides conventional knowledge editing benchmarks, there have also been some benchmarks designed for special use.
+Besides conventional knowledge editing datasets, there have also been some designed for special use.
 Mitchell et al.
-(2022b) propose ConvSent to edit LLMs’ sentiments on given topics.
-Ishibashi and Shimodaira (2024) propose Sanitation to forget specific information stored in LLMs and thus address privacy concerns.
+(2022b) propose ConvSent to edit LMs’ sentiments on given topics.
+Ishibashi and Shimodaira (2024) propose Sanitation to forget specific information stored in LMs and thus address privacy concerns.
 Huang et al.
 (2025) propose HalluEditBench to evaluate whether knowledge editing correct hallucinations.
-Our proposed CRAFT benchmark leverages statistical and financial data, which are both real-time and inherently dynamic.
+Our proposed CRAFT leverages statistical and financial data, which are both real-time and inherently dynamic.
 The strong interdependence among indicators allows each pair of edits to correspond to a meaningful reasoning portability test.
-This enables comprehensive evaluation across five dimensions: benchmark type, real-time property, real-world grounding, reasoning portability, and composite reasoning portability.
+This enables comprehensive evaluation across five dimensions: data type, real-time property, real-world grounding, reasoning portability, and composite reasoning portability.
 6.2 Knowledge Editing Methods
 Parameter-based Editing
-This line of work aims to edit model parameters or add extra parameters to perform knowledge editing.
+This line of work edits model parameters or adds extra parameters for knowledge editing.
 Meng et al.
-(2022)Meng et al.
+(2022), Meng et al.
 (2023) and Fang et al.
@@
 Mitchell et al.
-(2022a) adopts meta-learning Hospedales et al.
+(2022a) adopt meta-learning Hospedales et al.
 (2022) to predict the change of model parameters.
 Dong et al.
-(2022)Huang et al.
+(2022), Huang et al.
 (2023) and Wang et al.
@@
 Mitchell et al.
-(2022b) trains a separate counter-fact model and does not involve retrieval (thus categorized as parameter-based), failing to fully exploit the capability of LLMs themselves.
+(2022b) train a separate counter-fact model and does not involve retrieval.
 Retrieval-based Editing
@@
 Zheng et al.
-(2023) leverages in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting.
+(2023) leverage in-context examples of copying, updating and retaining knowledge to edit knowledge via prompting.
 Chen et al.
-(2024b) first prompts LLMs to decide the relevance of retrieved edits and then queries LLMs with or without in-context edits.
+(2024b) first prompt LMs to decide the relevance of retrieved edits and then query LMs with or without in-context edits, which may seem conceptually related to KEDAS but is essentially different in that it relies merely on prompting.
 Chen et al.
-(2024a) converts edits into soft prompt tokens with trained models and then concatenates user prompts to modify the behavior of frozen LLMs.
+(2024a) convert edits into soft prompt tokens with trained models and then concatenate user prompts to modify the behavior of frozen LMs.
 Jiang et al.
-(2024) aligns LLMs with knowledge editing by post-training with in-context edits.
+(2024) align LMs with knowledge editing by post-training with in-context edits.
 Our proposed KEDAS can also be categorized as retrieval-based.
 7 Conclusion
-In this work, we introduce CRAFT, a novel automatically updatable real-time benchmark for knowledge editing that focuses on national statistics and finance of China.
+We have introduced CRAFT, a novel automatically updatable real-time dataset for knowledge editing that focuses on national statistics and finance of China.
 We evaluate representative knowledge editing methods on CRAFT and reveal that existing methods have certain limitations.
-We then propose a simple yet effective approach, KEDAS, to better align LLMs with real-time knowledge editing.
+We then propose a simple yet effective approach, KEDAS, to better align LMs with real-time knowledge editing.
 Experiments validate the effectiveness of KEDAS on CRAFT in all aspects including edit success, locality and portability.
+We hope this work may serve as a catalyst for shifting the focus of knowledge editing from static update to dynamic evolution.
+Acknowledgments
+This work is supported by the National Natural Science Foundation of China (62076008).
 Limitations
-First, the CRAFT benchmark focuses on the Chinese language only (but our methodology is universal and can be applied to any other languages) and is limited to the domains of national statistics and finance (but it is easily expandable to domains that data are periodically updated and automatically accessible).
-Second, the composite reasoning portability of CRAFT is merely 2-hop (but can be expanded to multiple hops via combination of time and indicators).
-Third, the Finance subset of CRAFT can be updated annually only (but can be improved to be quarterly updated by substituting the annual financial statements to quarter ones).
-Fourth, only one Llama-3 LLM is employed due to limited time and a limited set of existing knowledge editing methods is evaluated in our experiment due to limited resources and compatibility issues.
+Domain
+CRAFT focuses on the Chinese language only and is limited to the domains of national statistics and finance.
+Nevertheless, our methodology is universal and can be applied to any other languages and CRAFT is easily expandable to domains that data are periodically updated and automatically accessible.
+Portability
+The composite reasoning portability of CRAFT is merely 2-hop. Nevertheless, it can be expanded to multiple hops via combination of time and indicators.
+Baselines
+Only a limited set of existing methods is evaluated in our experiments due to limited resources and compatibility issues.
+Filter
+The performance of KEDAS heavily depends on the filter. If the filter is not sufficiently robust, the overall performance of KEDAS will also be less competitive.
 Ethical Considerations
@@
 Reproducibility
-All the experiments are fully reproducible since all the methods are deterministic and sampling is disabled during LLM generation.
+Our work is fully reproducible since we have provided our implementation details and scripts for experiments.
 Potential Risks
@@
 BERT Apache
-paraphrase-multilingual-MiniLM-L12-v2 Apache
-EasyEdit MIT
-Llama-3 Llama 3 Community License Agreement
+MiniLM Apache
+EasyEdit & KnowEdit MIT
+Llama-3 Llama 3 Community License
+Qwen2.5 Apache
+Llama-2 Llama 2 Community License
 Table 8: Licenses of scientific artifacts we use.
@@
 We also include an out-of-scope input $q_{l}$ with the target $a_{l}$ that is without any retrieved edits.
-In this way, the LLM learns to leverage retrieved edits for relevant queries while keeping the answer unchanged for locality queries.
-Appendix B Details of LLM Alignment
+In this way, the LM learns to leverage retrieved edits for relevant queries while keeping the answer unchanged for locality queries.
+Appendix B Details of LM Alignment
 Hyperparameter Value
@@
 LoRA dropout 0
-Table 8: Hyper-parameters for LLM finetuning.
-We adopt LLaMA-Factory Zheng et al. (2024) to finetune the LLM. The alignment is done on an NVIDIA A40 48GB GPU and takes 2 hours.
+Table 9: Hyper-parameters for LM finetuning.
+We adopt LLaMA-Factory Zheng et al. (2024) to finetune the LM. The alignment is done on an NVIDIA A40 48GB GPU and takes 2 hours.
 The main training hyperparameters are presented in Table 9, most remain unchanged as the default values.
-Note that we train the LLM for only one epoch because finetuning for multiple epoches will cause over-fitting.
+Note that we train the LM for only one epoch because finetuning for multiple epoches will cause over-fitting.
+Prompt Template for Declarative You are a helpful assistant.
+You are given a query and a target new.
+Please generate the declaration form of the query and target new.
+Here is an example: Query: 2024年华大智造的总资产(亿元)是多少?
+Target new: 103.15 Declaration: 2024年华大智造的总资产(亿元)是103.15。 Please generate the declaration form of the query and target new below: Query: {query} Target new: {target_new} Declaration:
+Figure 2: Prompt template for the declarative form.
 Appendix C Details of Diverse Edit Augmentation
-Figure 2: Prompt template for the declarative form.
-Figure 3: Prompt template for the aliased form.
 The prompt templates of diverse edit augmentation for gpt-4o-mini are presented in Figure 2 and 3. We include manually written in-context examples in the templates to better instruct the model.
@@
 In this way, the filter learns to decide the relevance between the query and the edit.
-Appendix E Example from CRAFT-Statistics
+Appendix E Details Regarding KnowEdit
+Dataset ZsRE WikiBio $_{\text{recent}} WikiData$$_{\text{counterfact}}$
+# Train 10000 464 570 1428
+# Test 1304 306 1266 885
+Table 10: Data statistics of KnowEdit.
+Table 10 presents the data statistics of KnowEdit.
+Following Zhang et al.
+(2024) and Jiang et al.
+(2024), our evaluation metrics for KnowEdit include edit success (ES), locality (L) and portability (P).
+We also calculate the harmonic mean (HM) of the three scores as the edit quality.
+Note that WikiBio does not contain portability queries and we thus calculate the harmonic mean based on the first two metrics.
+Appendix F Other Baselines
+In our primary experiments, we also tried to evaluate many other modern baselines but encountered different sorts of issues (including compatibility, availability, cost, performance, etc.).
+[…]
+[…]
+MEND Mitchell et al.
+(2022a), SERAC Mitchell et al.
+(2022b) and RECIPE Chen et al.
+(2024a) require a lot of resources that we cannot afford currently for training and do not provide trained checkpoints for the Llama-3 herd of LMs used in our experiments.
+Therefore, we do not include these in our main experiments.
+Nevertheless, for a comprehensive comparison, we experiment on KnowEdit with all the mentioned baselines under the single editing setting using Llama-2-7B-Chat Touvron et al.
+(2023) due to that all the methods are compatible with single editing and the Llama-2 herd of LMs.
+The results are presented in Appendix I.
+Prompt Template for Aliased You are a helpful assistant.
+You are given a query, a target new, and a declaration.
+Please generate a paraphrased sentence of the declaration with the central term translated to English.
+Here is an example: Query: 2024年华大智造的总资产(亿元)是多少?
+Target new: 103.15 Declaration: 2024年华大智造的总资产(亿元)是103.15。 Paraphrased sentence: 2024年华大智造的Total Assets(100 million yuan)是103.15。 Here is another example: Query: 2024年9月中国的订销报纸份数当期值(万份)是多少?
+Target new: 137885.2 Declaration: 2024年9月中国的订销报纸份数当期值(万份)是137885.2。 Paraphrased sentence: 2024年9月中国的Issue of Newspapers, Current Period Value(10000 pieces)是137885.2。 Please generate a paraphrased sentence of the declaration below: Query: {query} Target new: {target_new} Declaration: {declaration} Paraphrased sentence:
+Figure 3: Prompt template for the aliased form.
+Appendix G Example from CRAFT-Statistics
   • Prompt 1: 2024年7月中国的货币和准货币(M2)供应量期末值(亿元)是多少?
@@
 Question: Which animal mentioned in the passage can predict storms?) Answer: C
-Appendix F Example from CRAFT-Finance
+Appendix H Example from CRAFT-Finance
   • Prompt 1: 2024年三友医疗的总资产(亿元)是多少? (English translation: What is the total assets (100 million yuan) of Sanyou Medical in 2024?) Target New: 23.07
@@
     • Prompt: 请根据以下材料回答问题。 (略)问题:这个故事告诉我们什么道理? (English translation: Read the following passage and answer the question. Question: What moral lesson does this story convey?) Answer: A
-Appendix G Results on Traditional Knowledge Editing Benchmarks
+Appendix I More Results on KnowEdit
+Model Method ZsRE WikiBio $_{\text{recent}} WikiData$$_{\text{counteract}}$
+ES L P HM ES L HM ES L P HM ES L P HM
+Llama-2 SERAC 94.9 44.9 40.6 52.2 98.4 70.1 81.9 92.0 33.6 47.0 48.4 78.0 29.3 42.7 42.6
+IKE 99.5 46.9 67.6 65.0 99.7 41.6 58.7 85.0 45.8 57.3 58.8 85.7 51.5 60.6 63.0
+EREN 36.7 67.1 47.8 47.6 60.8 64.0 62.4 44.4 59.8 38.5 46.0 25.6 67.8 25.4 32.2
+WISE 71.1 67.5 51.1 61.9 95.5 99.2 97.3 80.1 48.9 53.2 58.0 72.9 56.6 50.7 58.7
+RECIPE 96.4 60.0 52.6 65.1 63.8 61.9 62.8 80.8 61.2 51.4 62.2 55.6 66.8 36.8 49.9
+LTE 99.6 49.2 69.2 66.9 100.0 47.2 64.1 100.0 55.0 74.6 72.1 98.4 63.6 73.6 76.0
+KEDAS 99.6 91.0 73.6 86.7 100.0 100.0 100.0 100.0 77.6 75.9 83.1 98.4 76.3 77.1 82.8
+Qwen2.5 IKE 96.7 42.2 65.1 60.7 99.3 33.4 50.0 93.6 40.7 61.7 58.3 86.2 48.4 59.3 61.1
+EREN 34.4 60.7 46.0 44.6 58.1 53.0 55.4 42.1 55.6 37.4 43.8 21.5 62.4 22.0 27.8
+WISE 53.0 14.7 25.1 23.7 86.6 100.0 92.8 54.0 96.8 33.0 50.7 45.4 17.9 30.3 27.1
+LTE 99.1 48.0 65.4 64.9 98.4 38.7 55.5 99.7 42.5 68.9 62.4 98.0 49.4 69.9 67.0
+KEDAS 99.1 89.6 70.2 84.5 98.4 100.0 99.2 99.7 67.1 70.5 76.7 98.0 66.6 73.4 77.2
+Table 11: Main results of sequential editing. Metrics include edit success (ES), locality (L), portability (P) and harmonic mean (HM). The highest scores are bolded.
+The sequential editing results with Llama-2-7B-Chat Touvron et al.
+(2023) and Qwen2.5-7B-Instruct Qwen (2025) are presented in Table 11.
+KEDAS secures the highest harmonic mean scores on all the four datasets, outperforming previous methods significantly.
 Method ZsRE WikiBio $_{\text{recent}} WikiData$$_{\text{counteract}}$
 ES L P HM ES L HM ES L P HM ES L P HM
-IKE 98.5 43.5 64.4 61.7 96.9 37.9 54.5 97.4 48.8 67.2 65.7 91.3 53.7 61.4 65.4
-EREN 32.3 70.7 44.9 44.5 59.4 64.3 61.8 44.4 73.1 37.5 47.7 23.5 75.4 22.9 30.2
-WISE 60.0 51.3 40.4 49.3 84.5 95.9 89.8 68.8 94.5 43.1 62.1 47.1 45.4 35.0 41.8
-LTE 99.6 55.5 66.7 69.7 97.7 59.7 74.1 99.8 52.9 74.7 70.9 98.3 59.3 73.4 73.8
-KEDAS 99.6 90.5 71.4 85.5 97.7 100.0 98.9 99.8 73.9 76.1 81.7 98.3 69.2 77.4 79.9
-Table 9: Results on traditional knowledge editing benchmarks.
-Metrics include edit success (ES), locality (L), portability (P) and harmonic mean (HM) of the previous metrics.
-The highest scores are bolded.
-[…]
-The results are presented in Table 9.
-KEDAS secures the highest harmonic mean scores of edit success, locality and portability on all the four datasets, outperforming previous methods significantly.
-This further confirms that besides CRAFT, KEDAS is also effective on traditional knowledge editing benchmarks.
+LoRA 100.0 75.7 56.5 73.3 100.0 81.7 89.9 100.0 56.3 64.5 69.3 100.0 70.7 68.9 77.6
+MEND 97.9 96.4 61.1 81.2 58.3 90.8 71.0 94.4 77.9 43.3 64.5 - - - -
+ROME 96.7 53.7 52.7 62.5 96.0 62.7 75.8 97.0 54.7 55.7 64.5 98.5 50.4 55.8 62.6
+SERAC 97.0 60.0 61.3 69.3 98.4 72.8 83.7 99.0 36.4 64.9 56.6 99.0 32.4 76.7 55.6
+IKE 99.8 47.1 78.1 68.1 99.7 41.3 58.4 85.0 45.9 62.6 60.6 87.3 52.0 72.2 67.4
+EREN 36.7 67.1 47.8 47.6 60.8 64.0 62.4 44.4 59.8 38.5 46.0 25.6 67.8 25.4 32.2
+WISE 99.9 99.7 47.4 72.9 100.0 100.0 100.0 98.3 82.7 62.2 78.2 99.6 82.3 71.1 82.8
+AlphaEdit 97.5 66.9 51.2 67.0 99.0 71.8 83.2 97.8 60.8 52.3 65.5 98.9 64.1 47.7 64.3
+RECIPE 96.6 78.3 55.9 73.1 63.8 94.6 76.2 81.0 71.1 53.2 66.3 55.7 76.9 38.3 52.6
+LTE 99.9 49.2 77.6 69.4 100.0 47.2 64.1 100.0 55.0 82.0 74.3 100.0 65.0 88.0 81.6
+KEDAS 99.6 96.3 77.3 89.9 100.0 100.0 100.0 100.0 98.1 81.1 92.2 98.4 99.4 85.9 94.1
+Table 12: Results of single editing on KnowEdit with Llama-2-7B-Chat. The highest scores are bolded. “-” means the method encounters out-of-memory (OOM) or tensor overflow on our GPU.
+We also experiment with Llama-2-7B-Chat under the traditional single editing setting and the results are presented in Table 12.
+KEDAS exhibits the most balanced performance among all the 11 evaluated methods.